Total Pen Shipment Defective Percentage

A stationery store has decided to accept a large shipment of ball-point pens if an inspection of 20 randomly selected pens yields no more than two defective pens. Findthe probability that this shipment is accepted if 15% of the total shipment is defective.
Answer #1
x is bin(20,0.15)
P[shipment NOT accepted] = P[>2] = 1 - P[<=2] computed in the same way
= 1 - 0.4049 = 0.5951
answered by: keke